An x-ray has a wavelength of 4.40 Å. Calculate the energy (in J) of one photon of this radiation. Enter your answer in scientific notation.

1
Expert's answer
2021-03-15T09:29:42-0400

Let Energy in J be...... E(J)


"E(J)=h(j.s)\u00d7v(Hz)"


Where "v(Hz)=\frac{C(m/s)}{\lambda(m)}"


h = Planck's constant "6.63\u00d710^{-34}J.s"


C= Speed of light "3.00\u00d710^{8}m\/s"


Convert the wavelength from Angstrom Å to meter (m) "4.40 × \frac{1}{10^{10}}" = "4.40\u00d710^{-10}"


Therefore


"E(J) = 6.63×10^{-34}(\frac{3.00×10^8}{4.40×10^{-10}})"


"E(J) = 4.52\u00d710^{-16}"